# A new paradigm of mortality modeling via individual vitality dynamics

### v2

@@ Abstract
The significance of mortality modeling extends across multiple research areas, including life insurance valuation, longevity risk management, life-cycle hypothesis, and retirement income planning. Despite the variety of existing approaches, such as mortality laws and factor-based models, they often lack compatibility or fail to meet specific research needs. To address these shortcomings, this study introduces a novel approach centered on modeling the dynamics of individual vitality and defining mortality as the depletion of vitality level to zero. More specifically, we develop a four-component framework to analyze the initial value, trend, diffusion, and sudden changes in vitality level over an individual’s lifetime. We demonstrate the framework’s estimation and analytical capabilities in various settings and discuss its practical implications in actuarial problems and other research areas. The broad applicability and interpretability of our vitality-based modeling approach offer an enhanced paradigm for mortality modeling.
@@ 1 Introduction
Mortality modeling is a critical research topic, serving as the foundation of numerous research questions and applications in various fields. In actuarial science, accurate mortality models are essential for understanding and managing risks associated with life insurance products. They enable life actuaries to assess the underlying mortality risk of insurance policies and ensure the financial stability of life insurance companies. In the realm of long-term care and disability insurance, mortality models are also contributing as an essential tool, for example, in a multi-state setup to capture the complexities of transition probabilities and claim distributions.
@@
Beyond traditional insurance applications, mortality modeling plays an important role in retirement income planning as well. By estimating the survival time of individuals, mortality models can help to obtain optimal consumption and investment strategies, allowing for better financial planning and resource allocation throughout retirement. The ability to predict longevity and mortality trends is also crucial in pension fund management, public health planning, and demographic studies. In each of these areas, the reliability and interpretability of mortality modeling outcomes directly impact the quality of the decisions made.
@@
In the literature, a variety of mortality modeling approaches exist, each with its strengths and limitations. Mortality laws are among the oldest and most fundamental approaches. These laws, such as the Gompertz-Makeham law (Gompertz,, 1825; Makeham,, 1860), aim to describe age-specific mortality rates using simple mathematical functions. Mortality laws are widely appreciated for their simplicity and ability to capture the general trend of increasing mortality with age. Despite their historical significance and ease of use, these laws often fall short in capturing the complexities of real-life mortality data, especially in the presence of cohort effects and changing mortality trends over time. More recent studies on this approach include, for example, Gavrilov and Gavrilova, (2019), Li et al., 2021a , Di Palo, (2023) and Dang et al., (2023).
In contrast to traditional mortality laws, the Cox proportional hazards model (Cox,, 1972) represents a significant advancement by not assuming a specific baseline hazard function. This semi-parametric model allows the hazard function to vary over time, influenced by covariates such as demographic, health-related, or environmental factors. The flexibility of the Cox model in incorporating various risk factors makes it a powerful tool in survival analysis and epidemiology. However, one of its limitations is the proportional hazards assumption, which implies that the covariates have a multiplicative effect on the hazard function. This assumption may not hold in all cases, leading to potential biases in the estimated effects of covariates. A comprehensive review of the Cox model and its recent developments can be found in Kalbfleisch and Schaubel, (2023).
@@
Another category is developed from using continuous-time stochastic processes to describe the evolution of mortality intensities. The seminal work by Milevsky and Promislow, (2001) applied a continuous-time Cox model to survival analysis. Biffis, (2005) introduced the Markovian affine jump-diffusion process for modeling both financial and demographic risks. This model preserves analytical tractability and provides closed-form solutions to survival probabilities. Blackburn and Sherris, (2013) further validated the application of Markovian affine mortality models in longevity risk pricing and hedging. Bauer et al., (2010) focused on pricing longevity-linked securities using a forward force of mortality modeling approach. Other studies include, for example, Wong et al., (2017), Jevtić and Regis, (2019), Wang et al., (2021) and Zhou et al., (2022).
@@
Factor-based mortality models, leading by the seminal works of Lee and Carter, (1992) and Cairns et al., (2006), have become prominent due to their ability to capture underlying trends and volatility in mortality data. The Lee-Carter (LC) model and the Cairns-Blake-Dowd (CBD) model express mortality rates as a combination of age-specific components and time-varying factors to capture mortality changes over time. Significant advancements have been made in this direction, including the works of Haberman and Renshaw, (2012), Kleinow, (2015), Shang, (2019), and Hunt and Villegas, (2023), with a review provided by Cairns et al., (2009). Despite their popularity, these models are based on discrete-time settings, which may not be suitable for continuous-time research questions requiring explicit solutions.
@@
Similarly, spline-based mortality models offer a flexible non-parametric approach to mortality modeling by fitting smooth curves to mortality data. These models use splines, which are piece-wise polynomial functions, to capture the intricate patterns of mortality rates. Currie et al., (2004) demonstrated that the use of P-splines for smoothing mortality data provides a robust method for handling irregularities and fluctuations. Other studies include, for example, Dodd et al., (2018), Camarda, (2019) Hilton et al., (2019), Richards, (2020), Dodd et al., (2021), and Zhu and Zhou, (2023). The main advantage of spline-based models is their flexibility in fitting complex mortality patterns. However, their non-parametric nature can lead to overfitting and a lack of interpretability when compared to more structured models.
@@
Recently, an innovative approach of modeling mortality via an individual’s remaining survival energy has been proposed, distinguishing itself from traditional models that typically focus on populations or cohorts. Specifically, Shimizu et al., (2021) introduced the Survival Energy Model (SEM), suggesting that each individual is born with a certain initial level of survival energy that depletes stochastically over time through a time-inhomogeneous diffusion process. Building on this, Shimizu et al., (2023) expanded the SEM framework by incorporating more sophisticated stochastic processes to better capture the complexities of mortality dynamics. By focusing on the stochastic nature of an individual’s survival energy, the SEM provides a more intuitive understanding of mortality dynamics.
@@
We note that within the SEM framework, an individual’s death time is essentially the first time that the survival energy process reaches zero. Therefore, this modeling approach is closely connected to the field of first passage analysis, which is a well-established topic in applied probability with extensive applications in various fields such as operations management (see, e.g., Yamazaki, (2017) and Han et al., (2024)), quantitative finance (see, e.g., Davydov and Linetsky, (2001) and Cai and Kou, (2011)) and actuarial science (see, e.g., Lin and Willmot, (2000) and Li and Zhou, (2024)). We show in this paper that the depletion of an individual’s vitality (a.k.a., survival energy) exhibits similarities to the dynamics of an insurer’s cash flow. Specifically, the negative jumps in the vitality process represent decreases in vitality level due to accidents, analogous to how an insurer’s surplus process often takes a spectrally negative form to model claim payments.
@@
The concept of modeling individual vitality is not new beyond actuarial literature. It was pioneered by Strehler and Mildvan, (1960), who defined vitality as “the capacity of an individual organism to stay alive”. For a discussion on the application of the Strehler-Mildvan model to human mortality, please refer to Finkelstein, (2012) and the references therein. Anderson, (1992) further contributed by incorporating Brownian motion to construct a stochastic vitality process, transforming the study of mortality into a first-passage time problem. Numerous studies on mortality modeling have been conducted in this direction, including the works by Skiadas, (1995); Anderson, (2000); Weitz and Fraser, (2001); Aalen and Gjessing, (2004); Steinsaltz and Evans, (2004); Li and Anderson, (2009); Li et al., (2013); Li and Anderson, (2013); Skiadas and Skiadas, (2010, 2014); Li and Anderson, (2015); Sharrow and Anderson, (2016), and references therein.
The existing studies demonstrate that vitality modeling is a robust approach to understanding mortality dynamics. The recent consideration of survival energy models in actuarial science highlights their potential to offer new insights into mortality modeling and forecasting. By integrating the stochastic nature of individual vitality and considering both deterministic and stochastic components, vitality models can provide a comprehensive framework for mortality modeling that is both flexible and intuitive. The main objective of this paper is to show that vitality modeling can become a valuable tool in actuarial science, addressing a number of research problems and enhancing the solutions provided by existing methods. Moreover, vitality modeling offers an alternative, more intuitive understanding of mortality dynamics, which can complement and improve upon traditional actuarial approaches.
@@
The first contribution of this paper is the introduction of a vitality-based framework for mortality modeling. This framework conceptualizes mortality through the lens of individual vitality, where death is defined as the point at which an individual’s vitality depletes to zero. The fundamental idea is that each person is born with an initial level of vitality, which decreases over time due to both natural aging and external factors. The framework is built around four intuitive components: the initial vitality level, natural depletion trends, stochastic diffusion process, and sudden jumps. The resulting framework offers a comprehensive and flexible approach to mortality modeling and is capable of accommodating various features and ideas related to mortality dynamics. This flexibility sets the stage for our second contribution, where we demonstrate how different specifications of the proposed framework can lead to diverse and insightful modeling outcomes.
The second contribution of this paper lies in the various specifications of the proposed modeling framework, which is capable of representing existing mortality models, including the Gompertz law and the Cairns-Blake-Dowd model. By transforming these models into a unified vitality-based structure, we enhance their ability to capture more complex mortality dynamics, such as jumps to account for accidental deaths and diffusion to reflect the randomness in natural life decay. Under each specification, we derive survival probabilities and other relevant properties to showcase the insightful modeling outcomes, setting the stage for our third contribution on actuarial applications.
@@
The third contribution of this paper demonstrates the practical applications of the vitality-based modeling framework in actuarial science. We showcase that this framework can be employed in various actuarial tasks and has the potential to solve existing problems and improve traditional methods. Specifically, we discuss how vitality can be incorporated in calculating life expectancies, pricing life insurance products, and deriving optimal consumption decisions. Additionally, our vitality-based framework is utilized to model disability and recovery probabilities, analyze different causes-of-death, and explain the under- and over-estimation of survival probabilities throughout an individual’s life course.
@@
The structure of this paper is organized as follows. Section 2 details the proposed modeling framework, explaining its four key components: initial vitality level, natural depletion trends, stochastic diffusion processes, and sudden jumps. Section 3 covers the various specifications of the proposed framework, including static and dynamic cases, and offers specific examples to illustrate its estimation and analytical capabilities. Section 4 discusses the applications of the framework in actuarial science, such as life insurance valuation, disability modeling, and optimal consumption decisions. Section 5 presents an illustration on the numerical estimation of a vitality-based mortality model. Finally, Section 6 concludes the paper with a discussion on the limitations and suggestions for future research.
2 The Proposed Modeling Framework
In this section, we introduce the vitality-based mortality modeling framework, detailing its fundamental structure and components. The proposed framework conceptualizes mortality through the depletion of an individual’s vitality to zero and utilizes four components to describe the dynamics of the underlying vitality. Each component captures different aspects of vitality dynamics, jointly providing a comprehensive and flexible approach to mortality modeling.
@@ 2 The Proposed Modeling Framework
Consider an individual who is age $x$ at time 00 (the time of modeling), and let $V(t)$ be the vitality of this individual at time $t\geq 0$. The individual is considered deceased when the vitality level is exhausted (i.e., when $V(t)=0$). Let $\tau$ be the first passage time of the vitality $V(t)$ crossing zero; that is,
@@
The first component of our vitality-based modeling framework is the initial vitality of the individual being modeled, denoted as $V(0)$. This component is crucial as it sets the baseline from which all future vitality changes are measured. If the age $x$ is zero, then $V(0)$ represents the vitality level of a newborn. Otherwise, $V(0)$ can be interpreted as the remaining vitality level of the individual $x$ years after birth.
We can model $V(0)$ either as a fixed value or a random variable. A fixed value of $V(0)$ assumes that the level of initial vitality is known and also the same for all the individuals being modeled (i.e., all individuals who are age $x$ at time 00). This assumption might be reasonable if the cohort of individuals is homogeneous. For example, we can assume that all newborns from a certain birth year will have the exact same initial vitality. This simplification helps in creating a more streamlined model for analysis and is particularly useful in scenarios where there is little variation in initial health conditions within a cohort.
In a more realistic setting, the value of $V(0)$ should be treated as random to represent the inherent variability in the vitality level of individuals. This randomness reflects the diversity in individuals within a cohort, such as genetics, lifestyles, and other factors that affect an individual’s initial health condition. We should thus model $V(0)$ as a random variable with
@@
where $F_{0}$ is a distribution that could be, for example, exponential or Pareto. This approach acknowledges that even within a seemingly homogeneous cohort, individuals will have different initial vitality levels due to various reasons.
Furthermore, the remaining vitality level after $x$ years of living will differ among individuals due to their personal lifestyle, living environment, health status, and other individual-specific factors. Even for newborns, the initial vitality should vary within an inhomogeneous cohort to reflect different birth traits, such as genetic disorders, birth defects, and pregnancy conditions. This variability is essential for constructing a model that can accurately reflect individual differences, and also for building a linkage between existing mortality models and vitality-based models.
@@
The second component describes the decline in an individual’s vitality over time due to the natural aging process. We denote $Y(t)$ to represent the depletion trend in vitality that occurs as the individual ages beyond time 00. To reflect the general trend of vitality depletion in the absence of stochastic variations or sudden shocks, $Y(t)$ can be modeled as a deterministic function of time. In particular, we assume that
@@
where $\mu(t)$ is the depletion rate of vitality at time $t$ when the individual is aged $x+t$. The exact choice of $\mu(t)$ should be informed by both theoretical considerations and empirical studies.
The function $\mu(t)$ can be specified in various forms, depending on the specific characteristics and assumptions made. Common choices include constant, exponential, or more complicated functional forms that better fit empirical observations. For instance, an exponential decline might be used to represent a situation where the rate of vitality loss accelerates with age. Another example would be using the Gompertz law to specify $\mu(t)$; that is, $\mu_{x}(t)=bc^{x+t}$ where $b$ and $c$ are the conventional Gompertz parameters. We added a subscript of age $x$ to $\mu(t)$ to align with the typical notation of mortality laws and to emphasize that the depletion rate, given by the Gompertz law, depends on the individual’s age at time $t$.
@@
Acknowledging that an individual’s life is full of uncertainties and thus vitality should not be depleted in a perfectly predictable manner, the third component introduces a certain degree of uncertainty to the vitality dynamics of the individual. From the modeling perspective, given that the trend component $Y(t)$ exhibits a deterministic nature, it is reasonable to consider a diffusion component, denoted as $W(t)$, to reflect the fact that the vitality level of an individual can fluctuate randomly around its depletion trend. This stochastic variation will account for any minor health fluctuations that an individual may experience over time.
@@
where $B(t)$ is a standard Brownian motion and $\sigma(t)$ is a diffusion parameter at time $t$. A more advanced stochastic process can be considered to replace $B(t)$, such as a fractional Brownian motion, for capturing the long-term dependence in the individual’s health status over time. This setting is particularly useful when the individual has health conditions with a lingering effect on future vitality. Similar to the depletion rate, a piece-wise function of $W(t)$ can also be considered to reflect that different life periods may face different levels of vitality fluctuations.
With the diffusion component, we can simulate stochastic paths of individual vitality over time to provide insights into the variability and uncertainty inherent in the aging process. The simulations will generate a variety of potential future vitality trajectories, from which we can observe how different individuals’ path diverge from their deterministic trend. These simulated paths particularly help in understanding the probabilistic nature of vitality depletion and are crucial for applications such as deriving personalized optimal consumption decisions.
@@
Another natural phenomenon that could affect an individual’s health status is the occurrence of random events, such as an automobile incident or a medical surgery. These events can cause sudden and significant changes to an individual’s vitality level, but cannot be captured by the depletion trend or the stochastic diffusion. To address this, the last component of our proposed modeling framework is the jump component, denoted as $J(t)$. This component is designed to incorporate the impact of accidental events into the vitality dynamics via a jump process; that is,
@@
The jump component explains the variability in mortality due to accidental events that are not accounted for by gradual trends or random fluctuations. Together with the initial, trend and diffusion components, our proposed vitality-based framework is capable of providing insights into how a wide range of lifetime phenomena contribute to human mortality, including natural deaths, random health fluctuations, accidental events, and pre-existing medical conditions. By specifying these four components differently, our framework offers flexible modeling outcomes that encompass many existing mortality models. These different specifications and their implications will be discussed in the next section.

### v3

@@ Abstract
The significance of mortality modeling extends across multiple research areas, ranging from life insurance valuation to optimal lifetime decision-making. Existing approaches, such as mortality laws and factor-based models, often fall short in capturing the complexity of individual mortality, hindering their ability to address specific research needs. To overcome these limitations, this paper introduces a novel approach to mortality modeling centered on the dynamics of individual vitality. A four-component framework is developed to account for initial conditions, natural aging processes, stochastic fluctuations, and accidental events over an individual’s lifetime. We demonstrate the framework’s analytical capabilities across various settings and explore its practical implications in solving life insurance problems and deriving optimal lifetime decisions. Our results show that the proposed framework not only encompasses existing mortality models but also provides individualized mortality outcomes and offers an intuitive explanation for survival biases.
@@ 1 Introduction
Mortality modeling serves as a cornerstone for numerous research domains and practical applications, ranging from life insurance valuation and longevity risk management to optimal lifetime decision-making and public health policies. Accurate mortality models are essential for life insurers when evaluating and managing the demographic risks associated with life insurance products. These models are used to calculate prices, determine reserves, and fulfill regulatory requirements, ultimately facilitating better planning and management of life insurance products.
@@
Mortality modeling also plays an important role in optimal consumption and investment strategies. By estimating the survival time of individuals, mortality models can help develop optimal strategies for better financial planning and resource allocation. Additionally, the ability to predict longevity and mortality trends is crucial in pension fund management, public health planning, and demographic studies. In each of these areas, the reliability and interpretability of mortality modeling outcomes directly impact the quality of the decisions made.
Traditional mortality modeling approaches, such as mortality laws and factor-based models, have been instrumental in performing the aforementioned tasks. However, they often fall short in capturing the intricate dynamics of individual mortality, particularly when dealing with the heterogeneity in health conditions and accidental events over a lifetime. This gap in modeling capabilities highlights the need for a more individualized approach to better account for the variability and complexity of individual mortality.
This paper aims to address these limitations by introducing a novel mortality modeling framework centered on individual vitality dynamics. This framework conceptualizes mortality as the depletion of an individual’s vitality to zero and models the time of death as a first passage problem. By integrating multiple components into the underlying vitality process, our modeling approach delivers a more nuanced and intuitive understanding of mortality dynamics, which is important for improving actuarial practices and informing optimal decisions.
1.1 Literature review
@@
Mortality laws are among the oldest and most fundamental approaches in mortality modeling. These laws, such as the Gompertz-Makeham law (Gompertz,, 1825; Makeham,, 1860), aim to describe age-specific mortality rates using simple mathematical functions. Mortality laws are widely appreciated for their simplicity and ability to capture the general trend of increasing mortality with age, but they often fall short in capturing the complexities of real-life mortality data. More recent studies on this approach include, for example, Gavrilov and Gavrilova, (2019), Li et al., 2021a , and Di Palo, (2023).
@@
In contrast to mortality laws, another category is developed from using stochastic processes to describe mortality intensities. The seminal work by Milevsky and Promislow, (2001) applied a continuous-time Cox model to survival analysis. This modeling approach preserves analytical tractability and provides closed-form solutions to survival probabilities. Other studies include, for example, Biffis, (2005), Bauer et al., (2010), Blackburn and Sherris, (2013), Wong et al., (2017), and Zhou et al., (2022).
@@
Factor-based models, led by the seminal works of Lee and Carter, (1992) and Cairns et al., (2006), have become prominent due to their ability to capture the trend and volatility in real mortality data through a combination of age-specific components and time-varying factors. Significant advancements have been made in this direction, including the works of Haberman and Renshaw, (2012), Kleinow, (2015), and Hunt and Villegas, (2023), with a comprehensive review provided by Cairns et al., (2009).
@@
Similarly, spline-based models offer a non-parametric approach to mortality modeling by fitting smooth spline curves to capture the intricate patterns of mortality data. Currie et al., (2004) demonstrated that the use of P-splines for smoothing mortality data provides a robust method for handling irregularities and fluctuations. Other studies include, for example, Dodd et al., (2018), Camarda, (2019), Dodd et al., (2021), and Zhu and Zhou, (2023).
1.2 Motivation
@@
Recently, a new approach of modeling mortality has been proposed by Shimizu et al., (2021) and Shimizu et al., (2023). Specifically, Shimizu et al., (2021) introduces the Survival Energy Model (SEM), where each individual is assumed to have a survival energy level that depletes stochastically over time through a time-inhomogeneous diffusion process. Shimizu et al., (2023) expanded the SEM framework by using an inverse Gaussian process with an explicit mortality function. By focusing on the stochastic nature of an individual’s survival energy, the SEM provides an intuitive way of understanding mortality dynamics.
@@
Within the SEM framework, an individual’s death time is essentially the first passage time that the survival energy process crosses zero. This modeling approach is closely connected to the field of first passage analysis, a well-established topic in applied probability with extensive applications in operations management (see, e.g., Yamazaki,, 2017; Han et al.,, 2024), quantitative finance (see, e.g., Davydov and Linetsky,, 2001; Cai and Kou,, 2011), and actuarial science (see, e.g., Lin and Willmot,, 2000; Asmussen and Albrecher,, 2010). For instance, the depletion of an individual’s survival energy is similar to the dynamics of an insurer’s cash flow, where a negative jump in survival energy can be equivalently viewed as an insurer’s surplus process taking a spectrally negative form.
@@
The concept of modeling individual survival energy, also known as vitality, is not new beyond financial literature. It was pioneered by Strehler and Mildvan, (1960), who defined vitality as “the capacity of an individual organism to stay alive.” Anderson, (1992) further contributed by using a Brownian motion to describe vitality dynamics. Along this direction, numerous studies have been conducted, including the works by Anderson, (2000), Aalen and Gjessing, (2004), Li and Anderson, (2009), Skiadas and Skiadas, (2010), Sharrow and Anderson, (2016), and Anderson et al., (2017). A discussion on the application of the Strehler-Mildvan model to mortality modeling is available in Finkelstein, (2012).
1.3 Our contribution
The objective of this paper is to explore a novel approach to mortality modeling through the lens of individual vitality dynamics. By conceptualizing death as the depletion of an individual’s vitality to zero, we aim to derive mortality modeling outcomes that are individualized for unique health conditions, lifestyle factors, and random life events. We also attempt to apply the vitality-based modeling approach to a variety of emerging research problems. Specifically, this paper makes three significant contributions.
@@ 1.3 Our contribution
The first contribution of this paper is the introduction of a vitality-based framework for mortality modeling. This framework assumes that each individual has an initial vitality that depletes to zero over time due to both natural causes and external factors. It comprises four components: initial vitality level, natural depletion trends, stochastic diffusion, and sudden jumps. Our modeling approach offers a comprehensive and flexible method for capturing various aspects of mortality dynamics.
The second contribution lies in the various specifications of the proposed vitality-based framework. We demonstrate that certain specifications can represent existing mortality models, while others can generate more diverse modeling outcomes. Under each specification, we derive survival probabilities and other relevant properties to illustrate how the vitality-based framework can complement and improve existing models in describing complex mortality dynamics at the individual level.
@@
The third contribution of this paper relates to the practical applications of the vitality-based framework across several research domains. Specifically, we discuss how individual vitality can be incorporated into calculating life expectancies, pricing insurance products, and deriving optimal consumption decisions. Additionally, we derive disability and recovery probabilities, analyze different causes of death, and explain the under- and overestimation of survival probabilities under our vitality-based framework.
@@
The structure of this paper is organized as follows. Section 2 details the proposed vitality-based mortality modeling framework. Section 3 covers the various specifications of the proposed framework. Section 4 discusses how the vitality-based framework can be incorporated in various applications. Section 5 presents an illustration on the estimation of the proposed framework. Finally, Section 6 concludes the paper with a discussion on limitations and future research.
2 The Proposed Framework
@@
In this section, we introduce the vitality-based mortality modeling framework along with its fundamental structure and components. Consider an individual who is age $x$ at time 00 (the time of modeling), and let $V(t)$ be the vitality of this individual at time $t\geq 0$. The individual is considered deceased when the vitality level is exhausted (i.e., when $V(t)=0$). Let $\tau$ be the first passage time of the vitality $V(t)$ crossing zero; that is,
@@
The first component of our vitality-based modeling framework is the initial vitality of the individual being modeled, denoted as $V(0)$. This component sets the baseline from which all future vitality changes are measured. If the individual’s age $x$ is zero, then $V(0)$ represents the vitality level of a newborn. Otherwise, $V(0)$ can be interpreted as the remaining vitality level of the individual $x$ years after birth.
We can model $V(0)$ either as a fixed value or a random variable. A fixed value of $V(0)$ assumes that the level of initial vitality is known and also the same for all of the individuals being modeled. This setting may be reasonable if the group of individuals is homogeneous, for example, a portfolio of selected insurance policyholders. We may also assume that all newborns from a certain birth year will have the exact same initial vitality.
In a more realistic setting, the value of $V(0)$ should be treated as random to represent the inherent variability in the initial vitality level of individuals. This randomness reflects the diversity in individuals within a group, such as genetics, lifestyles, and other factors that affect an individual’s initial health condition. We would thus model $V(0)$ as a random variable with
@@
where $F_{0}$ is a distribution that could be, for example, exponential or Pareto.
Consider a group of individuals all aged $x>0$. Their remaining vitality after $x$ years of living will differ due to their personal lifestyle, living environment, and other individual-specific factors. Even for newborns with age $x=0$, the initial vitality should vary within a cohort year to reflect different birth traits, such as genetic disorders and birth defects. The random setting of $V(0)$ is useful for constructing a model that can reflect individual differences, and also for building a linkage between existing mortality models and vitality-based models.
@@
The second component describes the decline in an individual’s vitality due to the natural aging process. We denote $Y(t)$ to represent the depletion trend in vitality that occurs as the individual ages beyond time 00. First, $Y(t)$ can be modeled as a deterministic function of time. In particular, we can assume that
@@
where $\mu(t)$ is the vitality depletion rate at time $t$ when the individual is aged $x+t$. The exact choice of $\mu(t)$ should be informed by both theoretical considerations and empirical studies.
The function $\mu(t)$ can be specified in various forms, depending on the specific characteristics and assumptions made. Common choices include constant, exponential, or more complicated functional forms that better fit empirical observations. For instance, an exponential decline might be used to represent a situation where the rate of vitality loss accelerates with age. Another example would be using the Gompertz law to specify $\mu(t)$; that is, $\mu(t)=bc^{x+t}$ where $b$ and $c$ are the conventional Gompertz parameters.
@@
The third component introduces a certain degree of uncertainty to the vitality dynamics of the individual. From the modeling perspective, given that the trend component $Y(t)$ has a deterministic nature, it is reasonable to consider a diffusion component, denoted as $W(t)$, to reflect the fact that the vitality level of an individual can fluctuate randomly around its depletion trend. This stochastic variation will account for any vitality fluctuations that an individual may experience over time.
@@
where $B(t)$ is a standard Brownian motion and $\sigma(t)$ is a diffusion parameter at time $t$. A more advanced stochastic process can be considered to replace $B(t)$, such as a fractional Brownian motion for capturing the long-term dependence in the individual’s health status over time. Additionally, a piece-wise function of $W(t)$ can also be considered to reflect that different life periods may face different levels of vitality fluctuations.
With the diffusion component, we can simulate random trajectories of vitality over time to provide insights into the variability and uncertainty inherent in the aging process. These simulated trajectories are particularly helpful in understanding the probabilistic nature of vitality depletion and are crucial for applications such as deriving personalized optimal consumption decisions. Moreover, we can use simulated trajectories to analyze how different individuals’ vitality diverge from their deterministic trend.
@@
Another natural phenomenon that could affect an individual’s health status is the occurrence of random events, such as an automobile incident or a medical surgery. These events can cause sudden and significant changes to an individual’s vitality level, but cannot be captured by the depletion trend or the stochastic diffusion. To address this issue, we let the last component of our vitality-based modeling framework to be a jump component, denoted as $J(t)$.
To incorporate the impact of accidental events into the vitality dynamics, we apply a jump process to $J(t)$; that is,